solve the equation by completing the square x2+3x-9=0.

Answer by stanbon(75887)

solve the equation by completing the square
x^2+3x-9=0
---
x^2 + 3x + ? = 9 + ?
----
Complete the square:
x^2 + 3x + (3/2)^2 = 9 + (3/2)^2
---
Factor:
(x+3/2)^2 = (36/3)+(9/4) = 45/4
---
x + (3/2) + (9/2)sqrt(5) or x+(3/2) = (-9/2)sqrt(5)
----
x = (-3+9sqrt(5))/2 or x = (-3-9sqrt(5))/2
===============================
Cheers,
Stan H.